OpenStudy (anonymous):

Ming was planning a trip to Western Samoa. Before going, she did some research and learned that the exchange rate is 6 tala for $2. How many Tala would she get if she exchanged $6 ?

OpenStudy (tkhunny):

Add please. 6 tala for $2 12 tala for $4 etc.

OpenStudy (texaschic101):

you could make a proportion and solve

OpenStudy (texaschic101):

6/2 = x/6 -- (6 tala to $2 = x tala to $6) cross multiply and solve for x

OpenStudy (texaschic101):

or do it like tkhunny....that looks easier

OpenStudy (anonymous):

thank you very much even though i feel very stupid now but thank you.

OpenStudy (texaschic101):

don't feel stupid...believe me, I have asked questions on here and after I found out the answer...I felt the same way....but you shouldn't...we are just learning :)

OpenStudy (tkhunny):

And of course, never "cross-multiply"? There is no reason to multiply by 2 jsut to divide by it again. \(\dfrac{6}{2} = \dfrac{x}{6}\) Simplify the left \(3 = \dfrac{x}{6}\) Multiply by 6 \(18 = x\) Done.